# Mind the Privacy Unit! User-Level Differential Privacy for Language Model Fine-Tuning

### v2

@@ Abstract
Large language models (LLMs) have emerged as powerful tools for tackling complex tasks across diverse domains, but they also raise privacy concerns when fine-tuned on sensitive data due to potential memorization. While differential privacy ($\mathsf{DP}$) offers a promising solution by ensuring models are ‘almost indistinguishable’ with or without any particular privacy unit, current evaluations on LLMs mostly treat each example (text record) as the privacy unit. This leads to uneven user privacy guarantees when contributions per user vary. We therefore study user-level DP motivated by applications where it necessary to ensure uniform privacy protection across users. We present a systematic evaluation of user-level DP for LLM fine-tuning on natural language generation tasks. Focusing on two mechanisms for achieving user-level DP guarantees, Group Privacy and User-wise DP-SGD, we investigate design choices like data selection strategies and parameter tuning for the best privacy-utility tradeoff.
@@ 1 Introduction
Existing studies have explored $\mathsf{DP}$ fine-tuning techniques for various applications, such as natural language understanding (Li et al., 2022; Yu et al., 2021b; 2022), generation (Yu et al., 2022; Anil et al., 2022; Yue et al., 2023), and in-context learning (Tang et al., 2023; Wu et al., 2023; Duan et al., 2024). Many training algorithms and evaluations in the literature make the (implicit) assumption that each user only contributes a single record to the training dataset. This notion is usually referred to as Example-level DP or Record-level DP. In the context of language models, a record is a sequence of (tokenized) text. However, in many real-world scenarios, each user may provide multiple records [^1] For instance, patients contributing various medical records to train healthcare AI assistants, or users submitting different prompts/conversations to an open-domain chatbot.. Under Example-level DP, users contributing more records inevitably obtain weaker privacy protection — an undesirable outcome (see Figure 1). To ensure uniform privacy protection regardless of the number of records per user, we consider the stronger requirement of User-level DP, which ensures that each user obtains the same privacy guarantee. This shift from Example-level DP to User-level DP forms the core focus of our study.
@@ 2 User-level Differential Privacy Mechanisms
The definition of “neighboring datasets” is determined by the scope of privacy. For Example-level DP, two neighboring datasets are identical text corpora except for one text record. For User-level DP, we assume each user has a (variable) number of records, and two neighboring datasets are identical except for all the text belonging to a particular user.
@@
2:  $\sigma\leftarrow\textsc{PrivAccounting}(\frac{\varepsilon}{k},\frac{\delta(e^{\varepsilon/k}-1)}{e^{\varepsilon}-1},M,B,T)$ /* Stage 2. Prepare the dataset */
@@
Group Privacy. Thanks to the composition property of the DP guarantee, an Example-level DP algorithm can be directly turned into a User-level DP algorithm for the case where each user contributes at most $k$ records. Specifically, we can provide an upper bound on the privacy cost of Group Privacy by using the standard group privacy lemma.
Proposition 1 (Group Privacy; e.g., Vadhan (2017)).
If $\mathcal{M}$ satisfies $({\varepsilon},\delta)$-DP then for all $k\geq 1$, then the mechanism $\mathcal{M}^{\prime}$, that limits contribution from each user to be at most $k$ (in any arbitrary manner), satisfies $(k{\varepsilon},\delta\frac{e^{k{\varepsilon}-1}}{e^{\varepsilon}-1})$-user-level-DP.
Thus, if we choose $\sigma$ such that DP-SGD (with Example-level DP) would satisfy $({\varepsilon}^{\prime},\delta^{\prime})$-DP for ${\varepsilon}^{\prime}=\frac{{\varepsilon}}{k}$ and $\delta^{\prime}=\delta\frac{e^{{\varepsilon}^{\prime}}-1}{e^{{\varepsilon}}-1}$, then by Proposition 1, we have that Group Privacy DP-SGD satisfies $({\varepsilon},\delta)$-User-level DP.
@@
User-wise DP-SGD. Group Privacy can be applied to any underlying record-level DP algorithm. However, this flexibility comes at the potential cost of not being able to utilize the structure of the underlying algorithm directly. Here, we introduce a simple User-level DP algorithm, called User-wise DP-SGD, based on the classical DP-SGD algorithm (Abadi et al., 2016) for Example-level DP training. At each training step, User-wise DP-SGD samples a batch of users. For each user, it then samples several records, and computes the average gradients on those records [^3] The number of records sampled for each user does not need to be the same for different users. However, for ease of implementation, we keep them the same in our experiments.. The (average) gradients from the batch of users are then norm-clipped, mean-aggregated, and noised before passing to the optimizer. See Algorithm 2 for a formal algorithm description. We note the main modification of the original DP-SGD algorithm is user-wise sampling and clipping, which is similar to a combination of DP-SGD with FedAvg (McMahan et al., 2017) evaluated in some private federated learning tasks (McMahan et al., 2018; Kairouz et al., 2021). In this paper, we do not assume a federated learning setting, and allow the learner to inspect all user data to devise more sophisticated sampling strategies. Comparing to Group Privacy, User-wise DP-SGD does not need to preprocess the dataset to restrict that each user contributes at most $k$ records. And we do not even need to assume each user’s data consists of pre-defined records. For long form text, we can randomly sample contagious chunks during training, creating more diverse records.
@@
The algorithm in Asi & Liu (2024) (see Algorithm 3) is based on User-wise DP-SGD with a more involved gradient estimation sub-procedure. In the sub-procedure, the algorithm detects if the average gradients for the users in the batch are well-concentrated, in other words, if they are close enough to each other in $\ell_{2}$-distance. If most of the gradients are well-concentrated, the outliers are removed, and the remaining gradients are mean-aggregated and noised. It is shown that less noise is needed in this case. In the other case, if there is no strong concentration, the algorithm halts and refuses to output anything meaningful. Under the assumption that all the records are i.i.d. from some underlying distribution, this algorithm can provably achieve the optimal privacy-utility trade-off.
@@ 3 Experimental Setup
Our evaluation uses the Enron Email dataset (Klimt & Yang, 2004) and the BookSum dataset (Kryściński et al., 2022). We use the pretrained GPT-2 small (125M) model (Radford et al., 2019) and LoRA (Hu et al., 2022) fine-tuning, following Yu et al. (2021a). [^4] We note that the choice of DP fine-tuning recipe is orthogonal to the user-level DP algorithms, as the latter can be seamlessly combined with any DP fine-tuning pipeline. We show in Section C.2 that we observe consistent results with a different fine-tuning method (Li et al., 2022). Please see Section C.3 for results on larger GPT-2 models, and our analysis in Section C.1 showing that LoRA does not significantly impact model performance. For the Enron Email dataset, we treat each unique email sender as the privacy unit; for BookSum, we treat each book as the privacy unit. Example-level and privacy unit-level statistics for these two datasets are provided in Table 1.
There is no established consensus on the optimal method to sample records for training User-level DP mechanisms, whether for Group Privacy or User-wise DP-SGD. Therefore, in the following sections, we conduct a comprehensive evaluation exploring two key factors in this selection: 1) the record selection criteria, and 2) the number of records to select for each privacy unit. Please see Appendix B for the detailed experimental setup and Section C.4 for the analysis of the computational overhead.
@@ 4.2 Number of selected records ($k$) per unit
Random 35.28 31.50 28.51
Longest 35.00 31.38 28.29
Shortest 35.31 31.96 28.75
Highest PPL 35.30 31.58 28.46
Lowest PPL 35.28 31.69 28.59
@@
Random 27.34 27.17 27.02
Longest 27.38 27.17 27.05
Shortest 27.43 27.19 27.04
Highest PPL 27.42 27.19 27.06
Lowest PPL 27.39 27.16 27.03
@@
$\varepsilon=1.0$ 35.68 34.06 35.28
$\varepsilon=3.0$ 32.04 30.99 31.50
$\varepsilon=8.0$ 30.13 28.77 28.51
$\varepsilon=\infty$ 23.01 19.96 19.53
@@
$\varepsilon=1.0$ 27.31 27.34 27.45 27.56 27.49
$\varepsilon=3.0$ 27.14 27.15 27.17 27.24 29.68
$\varepsilon=8.0$ 27.08 27.05 27.02 27.07 27.22
@@
Optimal $k$ balances noise from privacy and training data diversity. To investigate the impact of $k$, we fix the data selection strategy to the random selection method and vary the value of $k$. We observe in Table 3 that for lower privacy budgets (i.e., $\varepsilon=1.0$ or $3.0$), using a smaller $k$ is generally better, as the undesirable increase in noise when $k$ increases outweighs the benefits of diverse data. However, for larger $\varepsilon$, using a larger $k$ is usually preferred, as the high privacy budget can potentially tolerate slightly more noise being introduced. In such cases, the diversity of the training data becomes more beneficial for larger $k$, as evidenced by lower perplexity for larger $k$’s when $\varepsilon=\infty$ (non-private setting) [^5] Note that for both methods, we report results for $\varepsilon=\infty$ following their corresponding algorithms in Algorithm 1 and Algorithm 2 while bypassing the stage of aggregation and noise addition.. We also observe similar results on GPT2 medium and large models, as detailed in Section C.3.
@@
5 User-wise DP-SGD
@@
5.1 Data selection strategy
5.2 Number of selected records ($k$) per unit
User-wise DP-SGD favors larger $k$, with diminishing returns.
In contrast to Group Privacy (Section 4.2), increasing $k$ monotonically improves User-wise DP-SGD’s utility by providing more diverse data per user. However, larger $k$ also incurs slightly higher computation (Section C.5). The benefit diminishes as $k$ grows, so for practical deployment, choose a sufficiently large $k$ to balance utility and efficiency. For instance, on the BookSum dataset, $k=10$ appears to achieve the best trade-off between these two factors.
5.3 Analysis of the advanced User-wise DP-SGD
As discussed in Section 2, Asi & Liu (2024) proposed an advanced User-wise DP-SGD algorithm that requires injecting less noise if per-user averaged gradients in a batch exhibit low variance. Specifically, recall that User-wise DP-SGD has a clipping norm parameter $C$, and the noise added is proportional to $C$. Asi & Liu (2024) showed that when most of the gradients are within $\ell_{2}$-distance $\tau$ of each other, adding noise proportional to $\tau$ suffices for privacy concern. Under the assumption that the records are drawn i.i.d. from some underlying distribution, Asi & Liu (2024) proved that this mechanism can achieve the optimal rate.
However, it remains unclear how applicable this (and many theoretically motivated) algorithm is to real world problems where the i.i.d. assumption might not be true, and the number of records per user is relatively small (e.g., $k\leq 50$). In order to answer this question, we measure the noise levels of standard User-wise DP-SGD and Asi & Liu (2024)’s method under different values of $\tau$, controlling the required gradient concentration. Figure 3 shows that for most $k$ and $\epsilon$, Asi & Liu (2024)’s noise is lower than User-wise DP-SGD’s only if $\tau/C<0.1$, making it highly unlikely to enter the advanced noising step (lines 1015 in Algorithm 3). Despite its potential benefits, this method relies on high user gradient concentration, which does not hold in our settings.
@@ 6 Comparing Group Privacy and User-wise DP-SGD
$\varepsilon=0.5$ 35.31 35.08
$\varepsilon=1.0$ 33.59 33.36
$\varepsilon=3.0$ 31.05 30.98
@@
User-wise DP-SGD outperforms Group Privacy. At the mechanism level, User-wise DP-SGD consistently outperforms Group Privacy, especially for smaller privacy budgets. This is potentially due to User-wise DP-SGD allowing for more diverse data being selected in different minibatches during training, while Group Privacy already constrains the training data to be a subset of user records during the preprocessing stage. Additionally, for some users who have fewer than $k$ records, Group Privacy will perform resampling to obtain $k$ records. This repetition may also negatively impact the model’s performance. However, it is worth noting that Group Privacy requires minimal changes to implement on top of existing record-level private training, whereas User-wise DP-SGD may require modifications to the data loading pipeline. In this regard, practitioners may need to consider the implementation difficulty when choosing between these two methods.
@@ 7 Related Work
Previous research has applied DP-SGD for language models (Ponomareva et al., 2023), primarily focusing on the record level. These applications include work in natural language understanding (Li et al., 2022; Yu et al., 2021b; 2022), natural language generation (Anil et al., 2022), in-context learning (Tang et al., 2023; Wu et al., 2023; Duan et al., 2024), prompt tuning (Li et al., 2023b), synthetic text generation (Al Aziz et al., 2021; Yue et al., 2023), and retrieval-augmented language models (Huang et al., 2023).

### v3

@@ Abstract
Large language models (LLMs) have emerged as powerful tools for tackling complex tasks across diverse domains, but they also raise privacy concerns when fine-tuned on sensitive data due to potential memorization. While differential privacy ($\mathsf{DP}$) offers a promising solution by ensuring models are “almost indistinguishable” with or without any particular privacy unit, current evaluations on LLMs mostly treat each example (text record) as the privacy unit. This leads to uneven user privacy guarantees when contributions per user vary. We therefore study user-level DP motivated by applications where it is necessary to ensure uniform privacy protection across users. We present a systematic evaluation of user-level DP for LLM fine-tuning on natural language generation tasks. Focusing on two mechanisms for achieving user-level DP guarantees, Group Privacy and User-wise DP-SGD, we investigate design choices like data selection strategies and parameter tuning for the best privacy-utility tradeoff.
@@ 1 Introduction
Existing studies have explored $\mathsf{DP}$ fine-tuning techniques for various applications, such as natural language understanding (Li et al., 2022; Yu et al., 2021; 2022), generation (Yu et al., 2022; Anil et al., 2022; Yue et al., 2023), and in-context learning (Tang et al., 2024; Wu et al., 2023; Duan et al., 2023). Many training algorithms and evaluations in the literature make the (implicit) assumption that each user only contributes a single record to the training dataset. This notion is usually referred to as Example-level DP or Record-level DP. In the context of language models, a record is a sequence of (tokenized) text. However, in many real-world scenarios, each user may provide multiple records [^1] For instance, patients contributing various medical records to train healthcare AI assistants, or users submitting different prompts/conversations to an open-domain chatbot.. Under Example-level DP, users contributing more records inevitably obtain weaker privacy protection — an undesirable outcome (see Figure 1). To ensure uniform privacy protection regardless of the number of records per user, we consider the stronger requirement of User-level DP, which ensures that each user obtains the same privacy guarantee. This shift from Example-level DP to User-level DP forms the core focus of our study.
@@
We note that an independent and concurrent work of Charles et al. (2024) also studies user-level DP for LLM fine-tuning. While their key observations are largely consistent with ours, we discuss some differences in Section 7.
@@ 2 User-level Differential Privacy Mechanisms
The definition of “neighboring datasets” is determined by the scope of privacy. For Example-level DP, two datasets are neighboring if one can be obtained from the other by adding/removing one text record. For User-level DP, we assume each user has a (variable) number of records, and two datasets are neighboring if one can be obtained from the other by adding/removing all the text belonging to a particular user.
@@
2:  $\sigma\leftarrow\textsc{GroupPrivAccounting}(k,\varepsilon,\delta,M,B,T)$ /* Stage 2. Prepare the dataset */
@@
[…]
@@
User-wise DP-SGD. While Group Privacy can be applied to any underlying record-level DP algorithm, this flexibility comes at the potential cost of not utilizing the structure of the underlying algorithm directly. Here, we introduce a simple User-level DP algorithm, called User-wise DP-SGD, based on the classical DP-SGD algorithm (Abadi et al., 2016) for Example-level DP training. At each training step, User-wise DP-SGD samples a batch of users. For each user, it then samples several records, and computes the average gradients on those records [^4] The number of records sampled for each user does not need to be the same for different users. However, for ease of implementation, we keep them the same in our experiments.. The (average) gradients from the batch of users are then norm-clipped, mean-aggregated, and noised before passing to the optimizer. See Algorithm 2 for a formal algorithm description. We note the main modification of the original DP-SGD algorithm is user-wise sampling and clipping, which is similar to a combination of DP-SGD with FedAvg (McMahan et al., 2017) evaluated in some private federated learning tasks (McMahan et al., 2018; Kairouz et al., 2021). In this paper, we do not assume a federated learning setting, and allow the learner to inspect all user data to devise more sophisticated sampling strategies. $\textsc{PrivAccounting}({\varepsilon},\delta,M,B,T)$ here refers to any method that returns a $\sigma$ such that DP-SGD with $M$ records, batch size $B$, run for $T$ iterations, satisfies $({\varepsilon},\delta)$-DP. Since Algorithm 2 treats “users” as “records”, it suffices to simply replace $M$ by $N$ (total number of users) and $B$ by $n$ (number of users sampled in a batch); see Section B.3 for more details. Comparing to Group Privacy, User-wise DP-SGD does not need to preprocess the dataset to restrict that each user contributes at most $k$ records. And we do not even need to assume each user’s data consists of pre-defined records. For long form text, we can randomly sample contiguous chunks during training, creating more diverse records.
@@
We also evaluate the algorithm in Asi & Liu (2024) (see Algorithm 3) that is based on User-wise DP-SGD with a more involved gradient estimation sub-procedure. In the sub-procedure, the algorithm detects if the average gradients for the users in the batch are well-concentrated, in other words, if they are close enough to each other in $\ell_{2}$-distance. If most of the gradients are well-concentrated, the outliers are removed, and the remaining gradients are mean-aggregated and noised. It is shown that less noise is needed in this case. In the other case, if there is no strong concentration, the algorithm halts and refuses to output anything meaningful. Under the assumption that all the records are i.i.d. from some underlying distribution, this algorithm can provably achieve the optimal privacy-utility trade-off.
@@ 3 Experimental Setup
Privacy units.
Models.
We use the pretrained GPT-2 small (125M) model (Radford et al., 2019) and LoRA (Hu et al., 2022) fine-tuning, following Yu et al. (2022). [^5] We note that the choice of DP fine-tuning recipe is orthogonal to the user-level DP algorithms, as the latter can be seamlessly combined with any DP fine-tuning pipeline. We show in Section C.2 that we observe consistent results with a different fine-tuning method (Li et al., 2022). Please see Section C.3 for results on larger GPT-2 models, and our analysis in Section C.1 showing that LoRA does not significantly impact model performance.
There is no established consensus on the optimal method for sampling records to train User-level DP mechanisms, whether using Group Privacy or User-wise DP-SGD. Therefore, in the following sections, we conduct a comprehensive evaluation exploring two key factors in this selection: (i) the record selection criteria, and (ii) the number of records to select for each privacy unit. Please see Appendix B for the detailed experimental setup and Section C.4 for the analysis of the computational overhead.
@@ 4.2 Number of selected records ($k$) per unit
Random 35.23 31.45 28.51
Longest 34.95 31.32 28.23
Shortest 35.24 31.94 28.70
Highest PPL 35.28 31.53 28.40
Lowest PPL 35.28 31.62 28.58
@@
Random 27.31 27.15 27.00
Longest 27.34 27.19 27.02
Shortest 27.40 27.18 27.03
Highest PPL 27.40 27.16 27.06
Lowest PPL 27.37 27.14 27.03
@@
$\varepsilon=1.0$ 35.67 34.26 35.23
$\varepsilon=3.0$ 32.01 30.92 31.45
$\varepsilon=8.0$ 30.13 28.74 28.51
$\varepsilon=\infty$ 23.00 19.88 19.53
@@
$\varepsilon=1.0$ 27.30 27.34 27.31 27.54 27.46
$\varepsilon=3.0$ 27.13 27.15 27.15 27.22 29.67
$\varepsilon=8.0$ 27.08 27.05 27.00 27.07 27.22
@@
Optimal $k$ balances noise from privacy and training data diversity. To investigate the impact of $k$, we fix the data selection strategy to the random selection method and vary the value of $k$. We observe in Table 3 that for lower privacy budgets (i.e., $\varepsilon=1.0$ or $3.0$), using a smaller $k$ is generally better, as the undesirable increase in noise when $k$ increases outweighs the benefits of diverse data. However, for larger $\varepsilon$, using a larger $k$ is usually preferred, as the high privacy budget can potentially tolerate slightly more noise being introduced. In such cases, the diversity of the training data becomes more beneficial for larger $k$, as evidenced by lower perplexity for larger $k$’s when $\varepsilon=\infty$ (non-private setting) [^6] Note that for both methods, we report results for $\varepsilon=\infty$ following their corresponding algorithms in Algorithm 1 and Algorithm 2 while bypassing the stage of aggregation and noise addition.. We also observe similar results on GPT-2 medium and large models, as detailed in Section C.3.
5 User-wise DP-SGD
5.1 Data selection strategy
@@
5.2 Number of selected records ($k$) per unit
User-wise DP-SGD favors larger $k$, with diminishing returns.
In contrast to Group Privacy (Section 4.2), increasing $k$ monotonically improves User-wise DP-SGD’s utility by providing more diverse data per user. However, larger $k$ also incurs slightly higher computation (Section C.5). The benefit diminishes as $k$ grows, so for practical deployment, choosing a sufficiently large $k$ will balance utility and efficiency. For instance, on the BookSum dataset, $k=10$ appears to achieve the best trade-off between these two factors.
@@ 5.2 Number of selected records ($k$) per unit
5.3 Analysis of the advanced User-wise DP-SGD
As discussed in Section 2, Asi & Liu (2024) proposed a User-wise DP-SGD algorithm that achieves the state-of-the-art excess rate and requires injecting less noise if per-user averaged gradients in a batch exhibit low variance. Specifically, recall that User-wise DP-SGD has a clipping norm parameter $C$, and the noise added is proportional to $C$. Asi & Liu (2024) showed that when most of the gradients are within $\ell_{2}$-distance $\tau$ of each other, adding noise proportional to $\tau$ suffices for privacy purposes. Under the assumption that the records are drawn i.i.d. from some underlying distribution, Asi & Liu (2024) proved that this mechanism can achieve the optimal rate.
However, it remains unclear how applicable this algorithm, and other such theoretically motivated ones, are to real-world problems where the i.i.d. assumption might not be true, and the number of records per user is relatively small (e.g., $k\leq 50$). In order to answer this question, we measure the noise levels of standard User-wise DP-SGD and Asi & Liu (2024)’s method under different values of $\tau$, controlling the required gradient concentration. Figure 3 shows that for most $k$ and $\epsilon$, Asi & Liu (2024)’s noise is lower than User-wise DP-SGD’s only if $\tau/C<0.1$, making it highly unlikely to enter the advanced noising step (lines 1015 in Algorithm 3). Despite its potential benefits, this method relies on high user gradient concentration, which does not hold in our settings.
@@ 6 Comparing Group Privacy and User-wise DP-SGD
$\varepsilon=0.5$ 35.28 35.08
$\varepsilon=1.0$ 33.50 33.36
$\varepsilon=3.0$ 31.01 30.98
@@
User-wise DP-SGD outperforms Group Privacy. At the mechanism level, User-wise DP-SGD consistently outperforms Group Privacy, especially for smaller privacy budgets. This is potentially due to User-wise DP-SGD allowing for more diverse data being selected in different minibatches during training, while Group Privacy already constrains the training data to be a subset of user records during the preprocessing stage. Additionally, for some users who have fewer than $k$ records, our implementation of Group Privacy will perform resampling to obtain $k$ records. This repetition may also negatively impact the model’s performance. We acknowledge this as a potential limitation of our evaluation that could be addressed in future work. However, it is worth noting that Group Privacy requires minimal changes to implement on top of existing record-level private training, whereas User-wise DP-SGD may require modifications to the data loading pipeline. In this regard, practitioners may need to consider the implementation difficulty when choosing between these two methods.
@@ 7 Related Work
Previous research has applied DP-SGD for language models (Ponomareva et al., 2023), primarily focusing on the record level. These applications include work in natural language understanding (Li et al., 2022; Yu et al., 2021; 2022), natural language generation (Anil et al., 2022), in-context learning (Tang et al., 2024; Wu et al., 2023; Duan et al., 2023), prompt tuning (Li et al., 2023b), synthetic text generation (Al Aziz et al., 2021; Yue et al., 2023), and retrieval-augmented language models (Huang et al., 2023).
@@
We note that an independent and concurrent work of Charles et al. (2024) also examines the empirical performance of user-level DP methods for fine-tuning LLMs. The key findings that User-wise DP-SGD usually outperforms Group Privacy in terms of utility are generally consistent between the two studies. Notably, their work provides valuable insights by offering an explicit methodology for selecting the group size in both mechanisms, whereas ours mainly performs ablations of various sizes. Moreover, our work differs from theirs in that we explore a broader design space for data selection strategies under various User-level DP mechanisms. In this expanded exploration, we observe substantial performance gaps between suboptimal and better data selection methods. In addition, their evaluation focuses on the larger StackOverflow and news datasets using a single 350M-parameter model while our work utilizes books and email datasets and considers models of various sizes. We also provide case studies for the User-level DP algorithm with the state-of-the-art excess rate.